OpenStudy (anonymous):

how to solve for s: ((0.154*s)-161700)/1800000=((0.154*s)-69300)/3000000

OpenStudy (anonymous):

Find the LCD. The LCD is 9,000,000 \[\frac{5(0.154s-161700)}{9000000}=\frac{3(0.154s-69300)}{9000000}\] Thus, \[5(.0154s-161700)=3(0.154s-69300)\] Distribute the 5 on the left side of the equation and the 3 on the right side. \[.77s-808500=.462s-207900\] Isolate s and solve: \[.308s=600600\] \[\frac{.308s}{.308}=\frac{600600}{.308}\] \[s=1950000\]